# Parallel Corpora for Machine Translation in Low-Resource Indic Languages: A Comprehensive Review

### v1

@@ 2.3 Multimodal Corpora
Multimodal corpora extend beyond traditional text-based datasets by incorporating multiple data types, such as text, speech, and visual information, into a unified dataset  [43]. These corpora are instrumental in building more comprehensive MT models capable of handling real-world scenarios involving multiple modalities  [44]. IndicMultiModal [45], for instance, provides text, speech, and image datasets aligned across multiple Indic languages, supporting research in multimodal translation, speech synthesis, and cross-lingual retrieval. Such corpora are particularly beneficial for applications in digital accessibility  [46], interactive AI assistants
@@ 2.3.1 Speech-to-Text corpora
Speech-to-text align spoken language with its textual translation, making them essential for speech translation and automatic speech recognition (ASR) systems  [47]. These datasets are particularly valuable for creating voice-enabled translation models and developing ASR systems for Indic languages  [48].
One of the important corpora in this category is CVIT-IIITH Mann ki Baat Corpus  [49],mined from Indian Prime Minister Narendra Modi’s Mann ki Baat speeches. Since these speeches are carefully prepared and delivered in formal Hindi with occasional English phrases, the dataset is well-suited for studying political speech translation and handling Hindi-English code-switching. However, given its highly structured nature, it may not fully capture the variability of spontaneous speech, which is often a challenge for ASR models. Compared to other datasets, this corpus is more domain-specific, focusing on political communication. A more extensive alternative is the PMIndia corpus, which, while not explicitly speech data, consists of transcriptions of spoken news content into text [50]. It extends beyond Mann ki Baat by providing spoken speech transcriptions with translations across multiple Indian languages. While both PMIndia and Mann ki Baat focus on government-related content, PMIndia includes a broader set of formal speeches, policy discussions, and governance-related material. This makes it valuable for multilingual speech translation systems, though, like Mann ki Baat, its primary limitation is that government discourse follows a standardized linguistic structure, lacking the variation seen in informal conversations or spontaneous speech.
Another dataset designed for a specific domain is the QED Corpus  [51]. It is a text data which is derived from the video transcripts. It focuses on educational video transcripts in English and Hindi. Which cover a wide range of topics, QED is optimized for academic discourse, including lecture-style content. This makes it particularly beneficial for ASR and translation models targeting online education platforms, academic lectures, and instructional content. QED ensures high-quality transcriptions and translations tailored for educational use cases. QED Corpus fills the gap in academic and instructional content, making it an essential resource for educational applications. The choice of corpus depends on the desired application—whether for structured government communication, spontaneous public discourse, or domain-specific speech processing.
@@ 2.4 Text-to-Speech Corpora
Two notable contributions in this area are AI4Bharat Indic-TTS  [52] and BhasaAnuvaad  [53], both of which provide extensive linguistic resources to improve speech synthesis and ASR systems. AI4Bharat Indic-TTS is an open-source TTS model that supports 13 Indic languages, including Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Punjabi, Sanskrit, Tamil, Telugu, and Urdu. It is designed to produce high-quality synthetic speech with various speaker styles, making it suitable for applications such as language learning, assistive technologies, and media content creation. The model provides fine-grained control over speech parameters, including pitch, speed, and voice modulation, ensuring natural and expressive speech synthesis. Complementing this, BhasaAnuvaad serves as a comprehensive dataset for speech translation, encompassing over 44,400 hours of speech and 17 million text segments across 13 Indic languages and English. By incorporating both mined and high-quality curated parallel speech data, BhasaAnuvaad is a valuable resource for developing ASR and TTS systems, as it enables sentence-to-audio alignment, facilitating the training of robust speech synthesis models. These initiatives significantly contribute to advancing TTS technology for Indic languages, fostering inclusivity and accessibility for a diverse user base.
@@ 2.4.1 Image-to-Text Corpora
Image-to-text corpora helps in advancing multimodal learning, enabling models to understand and generate text based on visual input  [54]. In the context of Indic languages, several high-quality datasets have been developed to support multimodal research, particularly in image captioning, visual question answering (VQA), and image-grounded translation  [55]. Among the most significant contributions are the Hindi Visual Genome, Bengali Visual Genome, and Malayalam Visual Genome datasets. The Hindi Visual Genome dataset  [56] contains 31K multimodal pairs aligned in Hindi-English, providing a rich resource for training models in tasks such as image captioning and cross-lingual understanding. Similarly, the Bengali Visual Genome  [57] offers 29K multimodal pairs in Bengali-English, while the Malayalam Visual Genome includes 29K multimodal pairs in Malayalam-English. These datasets are designed to bridge the gap in low-resource Indic languages for multimodal AI applications. They are particularly useful in training  [58] and evaluating MT models, cross-lingual retrieval systems, and vision-language models (VLMs)  [59] for Indic languages. These corpora contribute significantly to the development of multimodal AI for Indic languages, facilitating better captioning, improved VQA systems, and enhanced multilingual vision-language applications. By providing a strong benchmark for multimodal learning, they enable robust model training for real-world applications such as automated image description generation and visual assistive technologies in Indian languages.
@@ 2.5 Code-Switched Corpora
Code-switching, the practice of mixing two or more languages within a single conversation or sentence, is common in multilingual communities, including those using Indic languages  [60]. Code-switched corpora are crucial for developing translation models that can handle real-world conversations where users frequently switch between languages such as Hindi-English, Bengali-English, and Tamil-English  [61].
GLUECoS  [24] is a well known code-switched corpus, which contains Hindi-English and Bengali-English code-switched data and also datasets from social media platforms like Twitter and WhatsApp, where code-mixing is prevalent  [62]. Training MT systems on such corpora improves their ability to handle informal and conversational text  [63]. The PHINC (Parallel Hinglish Social Media Code-Mixed Corpus)  [64] consists of 13.7k Hinglish (Hindi-English) sentences, making it one of the most comprehensive resources for studying mixed-language usage in digital communication. It is particularly valuable for social media NLP tasks, where speakers often switch between Hindi and English within a single sentence.
Similarly, the IIIT-H en-hi-codemixed-corpus  [65] is a code-mixed dataset consisting of 6k English-Hindi sentences. Compared to PHINC, this corpus has a smaller dataset size but higher-quality annotation, ensuring accurate training data for models dealing with Hinglish content. Its focus on token-level annotations makes it especially useful for tasks such as word-level language identification and code-mixed text normalization. The CALCS 2021 Eng-Hinglish dataset  [66] provides 10k parallel sentence pairs, focusing on formal and informal contexts of Hinglish usage. Compared to PHINC and IIIT-H, CALCS is particularly valuable for machine translation between English and Hinglish, helping models bridge the gap between standard English and code-mixed vernacular speech.
@@ 3.1 Evaluation Metrics
Evaluating parallel corpora for Indic languages requires a combination of automatic evaluation metrics that compare machine-generated translations with human reference translations. Common metrics like BLEU  [67], METEOR  [68], COMET  [69], and Translation Edit Rate (TER)  [70] are frequently used, each offering different perspectives on translation quality. BLEU, which focuses on n-gram precision  [71], is a widely used metric but can struggle with Indic languages due to their lexical complexity and flexible word order. For instance, languages like Hindi and Tamil exhibit significant syntactic differences from English, which BLEU may fail to capture adequately. METEOR improves upon BLEU by incorporating recall, synonym matching, and stemming, which makes it more effective for languages with rich morphology and varied word forms, such as Hindi, Bengali, and Telugu  [72]. However, like BLEU, METEOR still struggles with capturing semantic meaning and context, which are crucial for languages with high syntactic divergence from English.
To address these limitations, COMET, a newer metric, utilizes neural embeddings and contextualized models to evaluate translations based on semantic similarity and contextual understanding  [73]. This makes COMET particularly valuable for Indic languages, where it is essential to capture contextual meaning and semantic equivalence rather than just surface-level n-gram matches  [69]. Despite its advantages, COMET requires pretrained models and significant computational resources, which may not always be feasible for resource-constrained settings  [74]. Additionally, TER measures the edit distance between the machine-generated translation and the reference  [75]. By counting the minimum number of insertions, deletions, or substitutions needed to transform one translation into the other, TER is particularly useful for identifying structural mismatches between languages, especially those with flexible sentence structures, such as Hindi and Tamil  [70]. However, TER focuses on structural alignment rather than semantic accuracy, making it a complementary metric rather than a stand-alone tool.
@@ 3.2 Human-Translated Evaluation Datasets
High-quality, human-translated evaluation datasets are essential for benchmarking machine translation models, ensuring that they are assessed on accurate and reliable translations  [76]. Unlike automatically mined corpora, these datasets are manually curated by professional translators, making them gold-standard resources for evaluating translation quality across different language pairs. Human-translated corpora are particularly crucial for low-resource languages, where the availability of clean, parallel data is often limited  [77].
@@
Several notable human-annotated evaluation datasets have been developed to facilitate rigorous benchmarking of multilingual machine translation systems. Among these, the FLORES-101 dataset  [78], developed by Meta AI, is one of the most comprehensive evaluation resources. It provides human-translated test sets for 101 languages, including 14 Indic languages, making it a critical benchmark for assessing translation models in diverse linguistic settings. Following its success, Meta AI expanded the dataset to FLORES-200  [78], covering 200 languages, including 24 Indic languages. FLORES-200 represents one of the largest human-annotated evaluation datasets for multilingual translation, allowing researchers to systematically analyze the performance of models across a wide range of linguistic families. Both FLORES-101 and FLORES-200 use n-way parallel translation, meaning each sentence is consistently translated across all supported languages, enabling direct multilingual comparisons. These datasets are highly useful for benchmarking, but due to their relatively small sentence count, they are not suited for large-scale training.
@@
Meta AI had also introduced No Language Left Behind (NLLB)  [79] a benchmarks for large-scale translation efforts, NLLB-Seed  [42], a small but valuable human-translated dataset specifically designed for evaluating very low-resource languages. This dataset focuses on five Indian languages—Kashmiri, Manipuri, Maithili, Bhojpuri, and Chhattisgarhi—where high-quality parallel data is scarce. While FLORES-200 provides extensive language coverage, it does not always include languages with very limited training data. NLLB-Seed  [80] fills this gap by prioritizing data quality and focusing on extremely low-resource languages, ensuring that translation models trained on scarce data sources can still be evaluated effectively. A complementary dataset, NLLB-MD (Multi-Domain)  [79], extends this effort by providing human-annotated parallel translations across three key domains: news, unscripted informal speech, and health. Unlike FLORES datasets, which contain mostly general-purpose text, NLLB-MD allows for more fine-grained evaluation of translation models across different content types, addressing challenges such as domain adaptation and stylistic variation in machine translation. This makes it a valuable resource for improving translation models that operate in specific fields such as journalism, healthcare, or conversational AI.
@@ 3.3 Domain Adaptation and Bias
The usability of parallel corpora for MT is contingent on their domain coverage and representativeness  [81]. Most of the Indic corpora like samantar, PMIndia exhibit a strong bias toward formal and government-regulated domains, including legislative proceedings, religious scriptures, and legal texts  [24]. While these datasets facilitate structured translation tasks, they lack the coverage necessary for informal and domain-specific language variations essential in social media, e-commerce, and medical translations. For example, Specialized corpora, such as TICO-19  [11], have been curated to enhance healthcare-related translation, significantly improving domain-specific MT performance. Expanding parallel corpora to include informal, code-switched datasets from platforms like Twitter, WhatsApp, and online forums is crucial for improving translation quality in conversational and low-resource settings. Beyond domain generalization, data size and language representation remain pivotal. While large-scale corpora such as Bhasha  [53] provide substantial bilingual sentence pairs, their distribution skews heavily toward high-resource languages like Hindi and Bengali, leaving low-resource languages such as Santali, Konkani, and Maithili significantly underrepresented  [82]. Addressing this imbalance requires techniques such as cross-lingual transfer learning, wherein models pretrained on high-resource Indic languages are fine-tuned on their low-resource counterparts  [83]. Gender bias, particularly in languages like bengali with grammatical gender agreement, often results in incorrect translations of gender-neutral references  [84]. Furthermore, political biases inherent in government-curated corpora such as PMIndia and Tico-19 may introduce ideological skew, influencing translation fidelity. Addressing such biases necessitates adversarial debiasing strategies, including counterfactual translation augmentation, reinforcement learning-based neutralization, and bias-aware adversarial training  [85]. Ensuring fair and inclusive translations mandates continuous evaluation and mitigation of systemic biases, particularly for Indic languages with diverse sociopolitical and linguistic landscapes.

### v2

@@ 2.3 Multimodal Corpora
Multimodal corpora extend beyond traditional text-based datasets by incorporating multiple data types, such as text, speech, and visual information, into a unified dataset  Baltrušaitis et al. (2019). These corpora are instrumental in building more comprehensive MT models capable of handling real-world scenarios involving multiple modalities  Li et al. (2020). IndicMultiModal Kothapalli et al. (2021), for instance, provides text, speech, and image datasets aligned across multiple Indic languages, supporting research in multimodal translation, speech synthesis, and cross-lingual retrieval. Such corpora are particularly beneficial for applications in digital accessibility  Sun et al. (2021), interactive AI assistants.
@@ 2.3.1 Speech-to-Text corpora
Speech-to-text align spoken language with its textual translation, making them essential for speech translation and automatic speech recognition (ASR) systems  Jouvet et al. (2019). These datasets are particularly valuable for creating voice-enabled translation models and developing ASR systems for Indic languages  Sitaram et al. (2020).
One of the important corpora in this category is CVIT-IIITH Mann ki Baat Corpus  Philip et al. (2021), mined from Indian Prime Minister Narendra Modi’s Mann ki Baat speeches. Since these speeches are carefully prepared and delivered in formal Hindi with occasional English phrases, the dataset is well-suited for studying political speech translation and handling Hindi-English code-switching. However, given its highly structured nature, it may not fully capture the variability of spontaneous speech, which is often a challenge for ASR models. Compared to other datasets, this corpus is more domain-specific, focusing on political communication. A more extensive alternative is the PMIndia corpus, which, while not explicitly speech data, consists of transcriptions of spoken news content into text Haddow and Kirefu (2020). It extends beyond Mann ki Baat by providing spoken speech transcriptions with translations across multiple Indian languages. While both PMIndia and Mann ki Baat focus on government-related content, PMIndia includes a broader set of formal speeches, policy discussions, and governance-related material. This makes it valuable for multilingual speech translation systems, though, like Mann ki Baat, its primary limitation is that government discourse follows a standardized linguistic structure, lacking the variation seen in informal conversations or spontaneous speech.
Another dataset designed for a specific domain is the QED Corpus  Lamm et al. (2021). It is a text data which is derived from the video transcripts. It focuses on educational video transcripts in English and Hindi. Which cover a wide range of topics, QED is optimized for academic discourse, including lecture-style content. This makes it particularly beneficial for ASR and translation models targeting online education platforms, academic lectures, and instructional content. QED ensures high-quality transcriptions and translations tailored for educational use cases. QED Corpus fills the gap in academic and instructional content, making it an essential resource for educational applications. The choice of corpus depends on the desired application—whether for structured government communication, spontaneous public discourse, or domain-specific speech processing.
@@ 2.4 Text-to-Speech Corpora
Two notable contributions in this area are AI4Bharat Indic-TTS  Kumar and S (2023) and BhasaAnuvaad  Jain et al. (2024), both of which provide extensive linguistic resources to improve speech synthesis and ASR systems. AI4Bharat Indic-TTS is an open-source TTS model that supports 13 Indic languages, including Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Punjabi, Sanskrit, Tamil, Telugu, and Urdu. It is designed to produce high-quality synthetic speech with various speaker styles, making it suitable for applications such as language learning, assistive technologies, and media content creation. The model provides fine-grained control over speech parameters, including pitch, speed, and voice modulation, ensuring natural and expressive speech synthesis. Complementing this, BhasaAnuvaad serves as a comprehensive dataset for speech translation, encompassing over 44,400 hours of speech and 17 million text segments across 13 Indic languages and English. By incorporating both mined and high-quality curated parallel speech data, BhasaAnuvaad is a valuable resource for developing ASR and TTS systems, as it enables sentence-to-audio alignment, facilitating the training of robust speech synthesis models. These initiatives significantly contribute to advancing TTS technology for Indic languages, fostering inclusivity and accessibility for a diverse user base.
@@ 2.4.1 Image-to-Text Corpora
Image-to-text corpora helps in advancing multimodal learning, enabling models to understand and generate text based on visual input  Guo et al. (2024). In the context of Indic languages, several high-quality datasets have been developed to support multimodal research, particularly in image captioning, visual question answering (VQA), and image-grounded translation  Özdemir and Akagündüz (2024). Among the most significant contributions are the Hindi Visual Genome, Bengali Visual Genome, and Malayalam Visual Genome datasets. The Hindi Visual Genome dataset  Parida et al. (2019) contains 31K multimodal pairs aligned in Hindi-English, providing a rich resource for training models in tasks such as image captioning and cross-lingual understanding. Similarly, the Bengali Visual Genome  Sen et al. (2021) offers 29K multimodal pairs in Bengali-English, while the Malayalam Visual Genome includes 29K multimodal pairs in Malayalam-English. These datasets are designed to bridge the gap in low-resource Indic languages for multimodal AI applications. They are particularly useful in training  Xue et al. (2024) and evaluating MT models, cross-lingual retrieval systems, and vision-language models (VLMs)  Bordes et al. (2024) for Indic languages. These corpora contribute significantly to the development of multimodal AI for Indic languages, facilitating better captioning, improved VQA systems, and enhanced multilingual vision-language applications. By providing a strong benchmark for multimodal learning, they enable robust model training for real-world applications such as automated image description generation and visual assistive technologies in Indian languages.
@@ 2.5 Code-Switched Corpora
Code-switching, the practice of mixing two or more languages within a single conversation or sentence, is common in multilingual communities, including those using Indic languages  Garg et al. (2021). Code-switched corpora are crucial for developing translation models that can handle real-world conversations where users frequently switch between languages such as Hindi-English, Bengali-English, and Tamil-English  Bali et al. (2014).
GLUECoS  Khanuja et al. (2020) is a well known code-switched corpus, which contains Hindi-English and Bengali-English code-switched data and also datasets from social media platforms like Twitter and WhatsApp, where code-mixing is prevalent  Sitaram et al. (2019). Training MT systems on such corpora improves their ability to handle informal and conversational text  Winata et al. (2021a). The PHINC (Parallel Hinglish Social Media Code-Mixed Corpus)  Srivastava and Singh (2020) consists of 13.7k Hinglish (Hindi-English) sentences, making it one of the most comprehensive resources for studying mixed-language usage in digital communication. It is particularly valuable for social media NLP tasks, where speakers often switch between Hindi and English within a single sentence.
Similarly, the IIIT-H en-hi-codemixed-corpus  Dhar et al. (2018) is a code-mixed dataset consisting of 6k English-Hindi sentences. Compared to PHINC, this corpus has a smaller dataset size but higher-quality annotation, ensuring accurate training data for models dealing with Hinglish content. Its focus on token-level annotations makes it especially useful for tasks such as word-level language identification and code-mixed text normalization. The CALCS 2021 Eng-Hinglish dataset  Appicharla et al. (2021) provides 10k parallel sentence pairs, focusing on formal and informal contexts of Hinglish usage. Compared to PHINC and IIIT-H, CALCS is particularly valuable for machine translation between English and Hinglish, helping models bridge the gap between standard English and code-mixed vernacular speech.
@@ 3.1 Evaluation Metrics
Evaluating parallel corpora for Indic languages requires a combination of automatic evaluation metrics that compare machine-generated translations with human reference translations. Common metrics like BLEU  Papineni et al. (2002), METEOR  Banerjee and Lavie (2005), COMET  Rei et al. (2020), and Translation Edit Rate (TER)  Snover et al. (2006) are frequently used, each offering different perspectives on translation quality. BLEU, which focuses on n-gram precision  Callison-Burch et al. (2006), is a widely used metric but can struggle with Indic languages due to their lexical complexity and flexible word order. For instance, languages like Hindi and Tamil exhibit significant syntactic differences from English, which BLEU may fail to capture adequately. METEOR improves upon BLEU by incorporating recall, synonym matching, and stemming, which makes it more effective for languages with rich morphology and varied word forms, such as Hindi, Bengali, and Telugu  Li et al. (2024). However, like BLEU, METEOR still struggles with capturing semantic meaning and context, which are crucial for languages with high syntactic divergence from English.
To address these limitations, COMET, a newer metric, utilizes neural embeddings and contextualized models to evaluate translations based on semantic similarity and contextual understanding  Sun and Wang (2024). This makes COMET particularly valuable for Indic languages, where it is essential to capture contextual meaning and semantic equivalence rather than just surface-level n-gram matches  Rei et al. (2020). Despite its advantages, COMET requires pretrained models and significant computational resources, which may not always be feasible for resource-constrained settings  Larionov et al. (2024). Additionally, TER measures the edit distance between the machine-generated translation and the reference  Stanchev et al. (2019). By counting the minimum number of insertions, deletions, or substitutions needed to transform one translation into the other, TER is particularly useful for identifying structural mismatches between languages, especially those with flexible sentence structures, such as Hindi and Tamil  Snover et al. (2006). However, TER focuses on structural alignment rather than semantic accuracy, making it a complementary metric rather than a stand-alone tool.
@@ 3.2 Human-Translated Evaluation Datasets
High-quality, human-translated evaluation datasets are essential for benchmarking machine translation models, ensuring that they are assessed on accurate and reliable translations  Yan et al. (2024). Unlike automatically mined corpora, these datasets are manually curated by professional translators, making them gold-standard resources for evaluating translation quality across different language pairs. Human-translated corpora are particularly crucial for low-resource languages, where the availability of clean, parallel data is often limited  Haddow et al. (2022).
@@
Several notable human-annotated evaluation datasets have been developed to facilitate rigorous benchmarking of multilingual machine translation systems. Among these, the FLORES-101 dataset  Guzmán et al. (2019), developed by Meta AI, is one of the most comprehensive evaluation resources. It provides human-translated test sets for 101 languages, including 14 Indic languages, making it a critical benchmark for assessing translation models in diverse linguistic settings. Following its success, Meta AI expanded the dataset to FLORES-200  Guzmán et al. (2019), covering 200 languages, including 24 Indic languages. FLORES-200 represents one of the largest human-annotated evaluation datasets for multilingual translation, allowing researchers to systematically analyze the performance of models across a wide range of linguistic families. Both FLORES-101 and FLORES-200 use n-way parallel translation, meaning each sentence is consistently translated across all supported languages, enabling direct multilingual comparisons. These datasets are highly useful for benchmarking, but due to their relatively small sentence count, they are not suited for large-scale training.
@@
Meta AI had also introduced No Language Left Behind (NLLB)  Team et al. (2022a) a benchmarks for large-scale translation efforts, NLLB-Seed  Team et al. (2022b), a small but valuable human-translated dataset specifically designed for evaluating very low-resource languages. This dataset focuses on five Indian languages—Kashmiri, Manipuri, Maithili, Bhojpuri, and Chhattisgarhi—where high-quality parallel data is scarce. While FLORES-200 provides extensive language coverage, it does not always include languages with very limited training data. NLLB-Seed  Team (2022) fills this gap by prioritizing data quality and focusing on extremely low-resource languages, ensuring that translation models trained on scarce data sources can still be evaluated effectively. A complementary dataset, NLLB-MD (Multi-Domain)  Team et al. (2022a), extends this effort by providing human-annotated parallel translations across three key domains: news, unscripted informal speech, and health. Unlike FLORES datasets, which contain mostly general-purpose text, NLLB-MD allows for more fine-grained evaluation of translation models across different content types, addressing challenges such as domain adaptation and stylistic variation in machine translation. This makes it a valuable resource for improving translation models that operate in specific fields such as journalism, healthcare, or conversational AI.
@@ 3.3 Domain Adaptation and Bias
The usability of parallel corpora for MT is contingent on their domain coverage and representativeness  Labaka et al. (2016). Most of the Indic corpora like samantar, PMIndia exhibit a strong bias toward formal and government-regulated domains, including legislative proceedings, religious scriptures, and legal texts  Khanuja et al. (2020). While these datasets facilitate structured translation tasks, they lack the coverage necessary for informal and domain-specific language variations essential in social media, e-commerce, and medical translations. For example, Specialized corpora, such as TICO-19  Anastasopoulos et al. (2020), have been curated to enhance healthcare-related translation, significantly improving domain-specific MT performance. Expanding parallel corpora to include informal, code-switched datasets from platforms like Twitter, WhatsApp, and online forums is crucial for improving translation quality in conversational and low-resource settings. Beyond domain generalization, data size and language representation remain pivotal. While large-scale corpora such as Bhasha  Jain et al. (2024) provide substantial bilingual sentence pairs, their distribution skews heavily toward high-resource languages like Hindi and Bengali, leaving low-resource languages such as Santali, Konkani, and Maithili significantly underrepresented  Resnik (1999). Addressing this imbalance requires techniques such as cross-lingual transfer learning, wherein models pretrained on high-resource Indic languages are fine-tuned on their low-resource counterparts  Lample and Conneau (2019). Gender bias, particularly in languages like bengali with grammatical gender agreement, often results in incorrect translations of gender-neutral references  Stanovsky et al. (2019). Furthermore, political biases inherent in government-curated corpora such as PMIndia and Tico-19 may introduce ideological skew, influencing translation fidelity. Addressing such biases necessitates adversarial debiasing strategies, including counterfactual translation augmentation, reinforcement learning-based neutralization, and bias-aware adversarial training  Sun et al. (2019). Ensuring fair and inclusive translations mandates continuous evaluation and mitigation of systemic biases, particularly for Indic languages with diverse sociopolitical and linguistic landscapes.